OpenStudy (anonymous):

PLEASE HELP!!! As well as finding the answer, could you explain how you'd solve this problem? Thanks! :) Michaela climbed up a mountain at an average speed of 3 mi/h. She climbed down at an average speed of 5 mi/h. What is Michaela’s average speed for the entire trip? Round to the nearest tenth.

OpenStudy (anonymous):

assuming that the distance up the mountain = the distance down the mountain, she covered 1/2 of the total distance at 3 mi/h and 1/2 of the total distance at 5 mi/h. Her average speed = 1/2 (3 mi/h + 5 mi/h) = 4 mi/h

OpenStudy (anonymous):

Suppose the mountain is 1 mile from bottom to top. So the trip up took 1/3 of an hour and the trip down took 1/5 of an hour. Average speed = total distance / total time = 2 mi / ( 1/3 h + 1/5 h ) = 3.75 mi /h.

OpenStudy (anonymous):

so the answer is 3.8